# Physically Feasible Semantic Segmentation

### v3

@@ Abstract
State-of-the-art semantic segmentation models are typically optimized in a data-driven fashion, minimizing solely per-pixel or per-segment classification objectives on their training data. This purely data-driven paradigm often leads to absurd segmentations, especially when the domain of input images is shifted from the one encountered during training. For instance, state-of-the-art models may assign the label “road to a segment that is located above a segment that is respectively labeled as “sky, although our knowledge of the physical world dictates that such a configuration is not feasible for images captured by forward-facing upright cameras. Our method, Physically Feasible Semantic Segmentation (PhyFea), first extracts explicit constraints that govern spatial class relations from the semantic segmentation training set at hand in an offline, data-driven fashion, and then enforces a morphological yet differentiable loss that penalizes violations of these constraints during training to promote prediction feasibility. PhyFea is a plug-and-play method and yields consistent and significant performance improvements over diverse state-of-the-art networks on which we implement it across the ADE20K, Cityscapes, and ACDC datasets. Code and models will be made publicly available.
@@ 1 Introduction
Figure 1: Left: A traffic light segment includes a building segment in a semantic annotation from the ground truth of the dataset, which is a feasible inclusion. Right: A bus segment includes a traffic sign segment (marked by red bounding boxes) in a semantic prediction, which is an infeasible inclusion. Our method addresses such physical infeasibilities in semantic segmentation. Best viewed on a screen and zoomed in.
@@
Semantic segmentation is a fundamental task in computer vision and enables many downstream applications. The last decade has witnessed dramatic advances in this dense prediction task, driven by the introduction of end-to-end learned network architectures for solving it [24, 27, 2, 3, 44, 4, 14, 36, 47, 6, 45] and the development of ever stronger backbones coming from image classification [32, 15, 37, 10, 23, 46]. Indeed, designing a suitable backbone for dense semantic prediction that balances global context aggregation with preservation of fine local details has been a primary research direction in semantic segmentation. However, merely optimizing the backbone architecture neglects the structured prediction character of the problem. In particular, the simple optimization of cross-entropy loss terms even by state-of-the-art, foundational network architectures [36, 6] essentially poses semantic segmentation as a set of independent pixel-level or—at best—segment-level prediction problems and only loosely promotes regularity of the complete image-level outputs via the shared computation in the network across neighboring pixels. Very few deep-learning-based works [17] have focused on designing objectives that originate from the interaction between different pixels in the output predictions, but even these objectives are formulated at a rather local level of pixel neighborhoods and not across large regions of the image. To the best of our knowledge, no previous method in semantic segmentation with deep learning has considered high-level pairwise objectives at the coarse level of semantic segments.
@@
Aiming at promoting conformity of semantic segmentation models to high-level constraints, we present a new method named Physically Feasible Semantic Segmentation, or PhyFea. Our method has been empirically motivated by the observation that various state-of-the-art semantic segmentation networks [41, 6, 45] make errors in their predictions which are not simple misclassifications but fundamentally violate physical constraints which are in force in images of real-world 3D scenes, as illustrated in Fig. 1, and which a human would easily avoid. In particular, based on their prior knowledge about the world, humans know that a traffic sign segment cannot be entirely included in/surrounded by a bus segment. We term such segment inclusions as infeasible inclusions and consider them as an exemplary type of physical constraint that can be exploited to provide a segmentation model with higher-level objectives based on the feasible spatial configurations of semantic segments in the output.
@@
While it could be possible to exhaustively consider all pairs of semantic classes in the segmentation dataset at hand and conceptually determine which of these pairs correspond to such infeasible inclusions, we argue that this knowledge can be directly extracted from the data. More specifically, our method first processes the training set offline before training in a single pass and automatically identifies all class pairs for which a segment of the one class includes a segment of the other at least once in the data. These are valid inclusion pairs (cf. Fig. 1-left) and PhyFea should and does not penalize similar inclusions in the outputs of the segmentation network. By contrast, pairs of co-occurring classes for which a complete inclusion of one class by the other is never found in the data are deemed as infeasible inclusion pairs (given the large size of the training set) and they contribute to a feasibility objective that we define for optimization during training.
@@
More specifically, the major contribution of PhyFea is a novel inclusion loss which is used as an additional objective besides standard cross-entropy for optimizing a generic segmentation network. Our inclusion loss is computed based on the softmax outputs of the employed network, similarly to cross-entropy, which allows its general application to the vast majority of deep-learning-based semantic segmentation models. The key idea of computing this loss is to identify “soft infeasible inclusions, which occur as regions in the softmax score maps of infeasible inclusion class pairs in which the ranking of the two softmax values is inverted compared to their surroundings. We recognize that this simply amounts to a grayscale morphological area opening [13] and derive a differentiable, parameter-free and thus efficient implementation of it, which is normally included in the forward and backward pass of standard backpropagation.
@@
We have achieved a consistent performance improvement with PhyFea across three central semantic segmentation benchmarks, i.e. ADE20K [49], Cityscapes [8], and ACDC [31], and across a variety of state-of-the-art semantic segmentation networks on which we have implemented our method, i.e. SegFormer [41], OCRNet [45], and Mask2Former [6]. Importantly, we have also observed a reduction of infeasible inclusions in our outputs.
@@ 2 Related Work
Figure 2: Overview of PhyFea. Top left: the complete network architecture, where the standard cross-entropy loss $l_{\text{ce}}$ from the baseline network $\phi$ is added to the inclusion loss $l_{\text{inclusion}}$ computed by PhyFea. Bottom: The pipeline of PhyFea. For the softmax outputs $P_{\phi}(I)$ of the network and for each pair of classes $(c_{i},c_{j})\in\mathcal{C^{\prime}}$ that signifies an infeasible inclusion where $c_{i}$ cannot include $c_{j}$, we take the difference of the respective softmax scores, $\text{P}_{\phi_{c_{j}}}(I)-\text{P}_{\phi_{c_{i}}}(I)$, and rectify it with a $\operatorname{ReLU}$. After concatenating all such $\mathcal{C^{\prime}}$ rectified difference maps channel-wise into $\hat{P}_{\phi}(I)$, we negate the latter and set its border pixels to 1. Then an iterative operation for area opening is performed $\mathbf{T}$ times on $\hat{P}_{\phi}(I)$. In each iteration, we perform max-pooling with a 3x3 kernel and a stride of 1, and the result is multiplied with $\hat{P}_{\phi}(I)$ element-wise. The final area-opened tensor channels differ from their counterparts in $\hat{P}_{\phi}(I)$ only across regions of incorrect inclusions. This element-wise difference is stored in $\mathbf{H}(I)$ and the $L_{1}$ norm of the latter, capturing both the spatial extent and the intensity of infeasible inclusions, constitutes the inclusion loss $l_{\text{inclusion}}$ employed in PhyFea.
Physical priors using local correlations.  Physical domain knowledge in the medical domain is implemented especially in MRA, PET, and MRI scan images concerning optimizing energy function or implementation in Markov Random Field [22] (a graphical model approach). Other approaches to implementing physical priors in the semantic segmentation model are AAF [17] (Adaptive Affinity Field) and a few similar approaches [14, 40, 9, 43, 50, 12, 1] which are very closely related to CRF (Conditional Random Field) [20]. In AAF, the predicted feature map of PSPNet [47] is compared with the input feature map using multiple-sized kernel windows using KL divergence as a loss function.
Physical priors via full supervision.  Zhang et al. [39] stated that physical priors can be implemented in a neural framework in four different ways. The first strategy is to integrate constraints into data by generating synthetic training datasets. The second strategy is to design non-trainable custom layers of physical operators and pre-conditioners in the DNN architecture to modify or shape feature maps calculated within the network to make them consistent with the prior knowledge. Implementing physical priors in semantic segmentation networks is also done by modern techniques using Self-attention [34, 48]. CCNet [16] is one such example and a few other similar approaches are [11, 38, 48, 50, 25]. In CCNet, ResNet-101 [46] is used as the backbone to extract the feature map and dilated convolution is applied to increase the spatial dimension of the feature map. Then using 1-D convolution to obtain the necessary parameters to apply self-attention map the semantic structures. In this model, the key factor is applying self-attention to capture the semantic relationships between the classes.
Physical priors in the loss function.  Physical priors can be enforced in the loss function as regularization loss as in the Potts model [33] in a discrete domain that penalizes appearance discontinuity between neighboring pixels. The regularization term formulated in the discrete setting is biased by the discrete grid and favors curves to orient along with the grid, e.g. in horizontal and vertical or diagonal directions in a 4-connected lattice of pixels. Previous attempts have been made to implement physical priors by incorporating shape information in the architecture [18, 42, 28, 21]. Shape information is a powerful semantic descriptor for specifying targeted objects in an image. Shape priors can be modeled in three ways: geometrical [35], statistical, and physical.
There are special types of neural architectures where physical priors can be implemented in loss functions. These types of networks are also called Physics Informed Neural Networks (PINNs) [29]. PINNs are differential equation solvers where a differential equation encodes the physical constraint. The neural network approximates the solution of the equation using boundary constraints. There are studies done in the geophysics domain to implement geophysical priors and laws in deep neural networks as regularization terms in loss functions for training the DNNs.
@@
Figure 3: Visualization of feature maps within PhyFea for a class pair corresponding to an infeasible inclusion. In this example semantic prediction of SegFormer [41] on Cityscapes [8], bus infeasibly includes traffic sign. In the blue frame, we first show the difference map between the softmax scores of the two classes using the coolwarm colormap, where blue tones indicate larger scores for traffic sign and red for bus, respectively. We rectify the difference between the two softmax maps and show only the respective channel of the 3D tensor $\hat{P}_{\phi}(I)$, where black indicates zeros and red tones indicate positive values. The area opening operation is performed on $\hat{P}_{\phi}(I)$ and the regions in which $\hat{P}_{\phi}(I)$ is positive (red) but which are not connected to the border, i.e. traffic sign segments infeasibly included in bus segments, are opened. Exactly these segments are isolated in the final tensor $\mathbf{H}(I)$ and are used to compute our inclusion loss $l_{\text{inclusion}}$.
Nosrati et al. [26] surveyed incorporating prior knowledge in medical image segmentation. This work reviews various ways to device prior knowledge information to improve image segmentation. The said physical priors are implemented with global or local optimization; as implementation of physical priors in continuous or discrete domains. Rother et al. [30] suggest ways to incorporate appearance priors in image segmentation using Gaussian mixture models [19]. In the literature, there are two ways to model the appearance: 1) learning the appearance in an adaptive manner during the segmentation procedure, and 2) knowing the appearance model before performing segmentation (e.g. by observing the appearance distribution of the training data). Comparison of our method with the related works.  All the above methods to enforce physical priors into a deep neural architecture do not take into account the semantic relation among the classes in the image. For example, in outdoor scenes, the sky should be above the road, a pole cannot be included by a bus or a sidewalk cannot be surrounded by the wall of a building. This shows that during training the model; is not being trained keeping in mind all these semantic relations among the classes. PINN although takes into account shape priors in the medical domain, it cannot extend semantic segmentation of general scenes, because we cannot enforce a fixed structure of any class in outdoor scenes unlike fixed shape anatomical structures, which are required to formulate the differential equation approximated in PINN.
3 PhyFea
@@ 3.1 Method Overview
Fig. 2 shows the architectural overview of PhyFea. A 2D semantic segmentation network $\phi(X)$ is fed with an image $\mathbf{I}\in\mathbb{R}^{(3\times H\times W)}$ and produces the initial output $\phi(\mathbf{I})\in\mathbb{R}^{(C\times H\times W)}$, where $C$ is the number of classes. PhyFea takes $\phi(\mathbf{I})$ as input and computes a novel inclusion loss, $l_{\text{inclusion}}$. We identify infeasible inclusions for class pairs of interest via a morphological area opening operation, which is fully differentiable and is normally included in the forward and backward pass during training. Our loss $l_{\text{inclusion}}$ is then added to the standard cross-entropy loss $l_{\text{ce}}$ for $\phi(X)$ to obtain the total loss as
$l_{\text{total}}=l_{\text{ce}}+\alpha{}l_{\text{inclusion}},0<\alpha<1.$ (1)
$l_{\text{total}}$ is normally backpropagated to optimize the weights of $\phi$. Thus, PhyFea is a differentiable morphological module that incorporates data-driven physical priors in the form of infeasible inclusions to improve the optimization and conformity of any segmentation network. Notably, PhyFea is fully parameter-free. In (1), $\alpha$ is a hyperparameter that is used to balance the inclusion loss and the standard cross-entropy loss.
Algorithm 1 Computing the set $\mathcal{U}$ of feasible inclusion class pairs from the training set ground-truth annotations $\mathcal{S}$. The set $\mathcal{P}$ contains all pairs of non-identical class labels for the dataset taxonomy. We aim to check whether the first element of the pair, $p$, can include the second one, $q$. We are taking 5 pixels as the minimum threshold for a contour of $q$ completely overlapping with any segment of $p$. Function C returns (i) the number of connected segments of class $q$ in the input annotation, and (ii) a discrete image, with each pixel labeled with its connected component ID or 0 for the background. Function F fills the holes of the mask of the connected component at hand. Then, function Cont returns a binary image in which only pixels belonging to the outer contour of the respective connected component are set. All variables except component are discrete images $\in\mathbb{N}^{(1\times H\times W)}$. Function NNZ returns the number of non-zero pixels of its input.
@@
4:$\mathcal{U}\leftarrow\emptyset$
5:for  $(p,q)\leftarrow\mathcal{P}$ do
6:     for  $I\leftarrow\mathcal{S}$ do
7:         mask_1 $\leftarrow\mathbbm{1}[I=p]\in\mathbb{R}^{1\times H\times W}$
8:         mask_2 $\leftarrow\mathbbm{1}[I=q]\in\mathbb{R}^{1\times H\times W}$
@@
10:         for  $l\leftarrow 1$ to noOFLabels do
11:              comp_mask $\leftarrow(\mathbbm{1}[\texttt{component}=l])$
@@
15:              if $\textbf{NNZ}(\texttt{result})=\textbf{NNZ}(\texttt{contour})\land\textbf{NNZ}(\texttt{contour})\geq 5$ then
16:                  $\mathcal{U}\leftarrow\mathcal{U}\cup\{(p,q)\}$
17:              end if
18:         end for
19:     end for
20:end for

### v4

# Optimizing against Infeasible Inclusions from Data for Semantic Segmentation through Morphology

@@ Abstract
State-of-the-art semantic segmentation models are typically optimized in a data-driven fashion, minimizing solely per-pixel or per-segment classification objectives on their training data. This purely data-driven paradigm often leads to absurd segmentations, especially when the domain of input images is shifted from the one encountered during training. For instance, state-of-the-art models may assign the label “road” to a segment that is included by another segment that is respectively labeled as “sky”. However, the ground truth of the existing dataset at hand dictates that such inclusion is not feasible. Our method, Infeasible Semantic Inclusions (InSeIn), first extracts explicit inclusion constraints that govern spatial class relations from the semantic segmentation training set at hand in an offline, data-driven fashion, and then enforces a morphological yet differentiable loss that penalizes violations of these constraints during training to promote prediction feasibility. InSeIn is a light-weight plug-and-play method, constitutes a novel step towards minimizing infeasible semantic inclusions in the predictions of learned segmentation models, and yields consistent and significant performance improvements over diverse state-of-the-art networks across the ADE20K, Cityscapes, and ACDC datasets. https://github.com/SHAMIK-97/PhyFea_on_Mask2Former/tree/main.
@@ 1 Introduction
Figure 1: Left: A traffic light segment includes a building segment in a semantic annotation from the ground truth of the dataset, which is a feasible inclusion. Right: A bus segment includes traffic sign segments (in the red boxes) in a semantic prediction, which is an infeasible inclusion. Our method addresses such physical infeasibilities in semantic segmentation. Best viewed on a screen and zoomed in.
@@
Semantic segmentation is a fundamental task in computer vision and enables many downstream applications. The last decade has witnessed dramatic advances in this dense prediction task, driven by the introduction of end-to-end learned network architectures for solving it [30, 34, 5, 6, 47, 7, 20, 41, 50, 9, 48] and the development of ever stronger backbones coming from image classification [37, 21, 42, 14, 29, 49]. Indeed, designing a suitable backbone for dense semantic prediction that balances global context aggregation with preservation of fine local details has been a primary research direction in semantic segmentation. However, merely optimizing the backbone architecture neglects the structured prediction character of the problem. In particular, the simple optimization of cross-entropy loss terms even by state-of-the-art network architectures [41, 9] essentially poses semantic segmentation as a set of independent pixel-level or- at best—segment-level prediction problems and only loosely promotes regularity of the complete image-level outputs via the shared computation in the network across neighboring pixels. Very few deep-learning-based works [26] have focused on designing objectives that originate from the interaction between different pixels in the output predictions, and those that do focus only on local neighborhoods rather than large image regions. To the best of our knowledge, no previous method in semantic segmentation with deep learning has considered high-level pairwise objectives at the coarse level of semantic segments.
@@
Aiming at promoting conformity of semantic segmentation models to high-level constraints, yet adhering to the taxonomy of the dataset at hand, we present a new method named Infeasible Semantic Inclusions (InSeIn). Our method has been empirically motivated by the observation that various state-of-the-art semantic segmentation networks [44, 9, 48, 1] make errors in their predictions, which are misclassifications or false positives, as illustrated in Fig. 1. In particular, based on the Cityscapes [11] taxonomy, a road segment cannot be entirely included in/surrounded by a sky segment, and the same holds for traffic sign and bus respectively. We term such segment inclusions infeasible inclusions and use them as an exemplary type of constraint that can provide a segmentation model with higher-level conformity based on the feasible spatial configurations of semantic segments in the ground truth.
@@
While it is possible to exhaustively consider all pairs of semantic classes in the dataset at hand and manually determine which of them correspond to such infeasible inclusions, we argue and experimentally evidence that infeasible inclusions can be directly identified from large-scale data. More specifically, InSeIn first processes the training set offline before training in a single pass and automatically identifies all class pairs for which a segment of the one class includes a segment of the other at least once in the data. InSeIn prunes this inclusion pair set by thresholding the relative frequency of each inclusion with respect to the co-occurrence of the respective pair to limit the impact of potential noise in the labels. The pairs obtained after pruning are valid inclusion pairs (cf. Fig. 1-left), and InSeIn should and does not penalize such inclusions in the outputs of the segmentation network. By contrast, pairs of co-occurring classes for which a complete inclusion of one class by the other is never found in the large-scale data are deemed as infeasible inclusion pairs, and they contribute to a feasibility objective that we define for optimization during training.
@@
More specifically, the major contribution of InSeIn is a novel inclusion loss, which is used as an additional objective besides standard cross-entropy for optimizing a generic segmentation network. Our inclusion loss is computed based on the softmax outputs of the employed network, similarly to cross-entropy, which allows its general application to the vast majority of deep-learning-based semantic segmentation models. The key idea of computing this loss is to identify “soft” infeasible inclusions, which occur as regions in the softmax score maps of infeasible inclusion class pairs in which the ranking of the two softmax values is inverted compared to their surroundings. We recognize that this simply amounts to a grayscale morphological area opening [16] and derive a differentiable, parameter-free, and efficient implementation of it, which is normally included in the forward and backward pass of standard backpropagation.
@@
We have achieved a consistent performance improvement with InSeIn across three central semantic segmentation benchmarks, i.e. ADE20K [52], Cityscapes [11], and ACDC [36], and across a variety of state-of-the-art semantic segmentation networks on which we have implemented our method, i.e. SegFormer [44], OCRNet [48], Mask2Former [9], OneFormer [25] and SegMAN [15]. Importantly, we have observed a substantial quantitative reduction of infeasible inclusions in our semantic predictions.
@@ 2 Related Work
Figure 2: Overview of InSeIn. Top left: the complete network architecture, where the standard cross-entropy loss $l_{\text{ce}}$ from the baseline network $\phi$ is added to the inclusion loss $l_{\text{inclusion}}$ computed by InSeIn. Bottom: The pipeline of InSeIn. For the softmax outputs $P_{\phi}(I)$ of the network and for each pair of classes $(c_{i},c_{j})\in\mathcal{C^{\prime}}$ that signifies an infeasible inclusion where $c_{i}$ cannot include $c_{j}$, we take the difference of the respective softmax scores, $\text{P}_{\phi_{c_{j}}}(I)-\text{P}_{\phi_{c_{i}}}(I)$, and rectify it with a $\operatorname{ReLU}$. After concatenating all such $\mathcal{C^{\prime}}$ rectified difference maps channel-wise into $\hat{P}_{\phi}(I)$, we negate the latter and set its border pixels to 1. After that, an iterative operation for area opening is performed $T$ times on $\hat{P}_{\phi}(I)$. In each iteration, we perform max-pooling with a 3x3 kernel and a stride of 1, and the result is multiplied with $\hat{P}_{\phi}(I)$ element-wise. The final area-opened tensor channels differ from their counterparts in $\hat{P}_{\phi}(I)$ only across regions of incorrect inclusions. This element-wise difference is stored in $\mathbf{H}(I)$ and the $L_{1}$ norm of the latter, capturing both the spatial extent and the intensity of infeasible inclusions, constitutes the inclusion loss $l_{\text{inclusion}}$ employed in InSeIn.
Semantic Segmentation.  Semantic segmentation deals with pixel classification. Pre-deep learning era methods mainly used lines and contour detection to classify similar pixel groups. Pixel misclassifications are mainly associated with boundary regions of any class segment and are studied in [31]. Here they classified pixel errors into 3 categories: false positives where the predicted edges erroneously appear in non-boundary areas, merging error occurs when the predicted results fail to capture the corresponding edges, resulting in false negatives and displacement error arises when the predicted edges deviate from their true positions. Such displacements may occur when the image features are misaligned. Pre-deep learning era attempted to segment misclassified pixels using graph searching, Laplace operator and other heuristic methods [39, 32, 45, 17, 18, 35].
Modern semantic segmentation methods.  The challenge to classify pixels in semantic segmentation is studied in [27] where the authors designed an architecture comprised of several sub-models arranged in a layer cascade fashion. Another paper [12, 46, 43] where the authors devised a framework for nighttime segmentation, where the input data is highly noisy. In this work, they have designed a seperate module to segment hard regions in an image and another module to segment the easy regions. Finally they merged the output of 2 modules to get the final prediction. Another paper  [13] where authors divided the input image into several patches and visualized the whole image as a unidirectional acyclic graph where the vertices are patches. Then they constituted a sperate ground truth which constitutes a seperate class for the boundaries and optimzed the model with the newly formed ground truth.
Contribution of loss function in Semantic segmentation.  Pixel misclassification problem in segmentation task is handled using depth cues in semantic segmentation in the paper [19, 33]. Here the authors constituted 2 loss-weight modules which output a loss weight map by employing two depth-related measurements of hard pixels: Depth Prediction Error and Depth-aware Segmentation Error. The loss weight map is optimized with segmentation loss to train the network. Another paper [4] devised a custom loss function to tackle misclassification by leveraging the sigmoid function to increase the loss and loss derivative of misclassified and uncertain pixels. Another popular method is Binary Focal Loss [28, 2, 24, 23, 40] where the author used the normal cross entropy loss and down-weighted the easily classified pixels. Another paper [38] developed a region wise(RW) loss through pixel-wise multiplication of RW map and softmax output for biomedical image segmentation.
All the above methods attempted to solve false positive regions at the pixel level. None of them takes into account the pairwise semantic relationship between the classes, as we do in our approach. We have also observed that false positives occur not only at the hard regions but also in easy-to-classify segments. We focus on the case of false positive segments in the form of infeasible inclusions and aim at optimizing against this type of error.
@@
Figure 3: Visualization of feature maps within InSeIn for a class pair corresponding to an infeasible inclusion. In this example semantic prediction of SegFormer [44] on Cityscapes [11], bus infeasibly includes traffic sign. In the blue frame, we first show the difference map between the softmax scores of the two classes using the coolwarm colormap, where blue tones indicate larger scores for traffic sign and red for bus, respectively. We rectify the difference between the two softmax maps and show only the respective channel of the 3D tensor $\hat{P}_{\phi}(I)$, where black indicates zeros and red tones indicate positive values. The area opening operation is performed on $\hat{P}_{\phi}(I)$ and the regions in which $\hat{P}_{\phi}(I)$ is positive (red) but which are not connected to the border, i.e. traffic sign segments infeasibly included in bus segments, are opened. Exactly these segments are isolated in the final tensor $\mathbf{H}(\mathbf{I})$ and are used to compute our inclusion loss $l_{\text{inclusion}}$.
3 InSeIn
@@ 3.1 Method Overview
Fig. 2 shows the architectural overview of InSeIn. A 2D semantic segmentation network $\phi(X)$ is fed with an image $\mathbf{I}\in\mathbb{R}^{3\times H\times W}$ and produces the initial output $\phi(\mathbf{I})\in\mathbb{R}^{C\times H\times W}$, where $C$ is the number of classes. InSeIn takes $\phi(\mathbf{I})$ as input and computes a novel inclusion loss, $l_{\text{inclusion}}$. We identify infeasible inclusions for class pairs of interest via a morphological area opening operation, which is fully differentiable and is normally included in the forward and backward pass during training. Our loss $l_{\text{inclusion}}$ is then added to the standard cross-entropy loss $l_{\text{ce}}$ for $\phi(\mathbf{I})$ to obtain the total loss as
$l_{\text{total}}=l_{\text{ce}}+\alpha{}l_{\text{inclusion}},0<\alpha<1.\vspace{-2mm}$ (1)
$l_{\text{total}}$ is normally backpropagated to optimize the weights of $\phi$. Thus, InSeIn is a differentiable morphological module that enhances the optimization and conformity of any segmentation network by integrating high-level data-driven constraints in the form of infeasible inclusions. Notably, the InSeIn module is both learned parameter-free and differentiable, allowing gradients to propagate through the network during training. The only extra parameter of InSeIn is $\alpha$ in (1), a hyperparameter which is used to balance the inclusion loss and the standard cross-entropy loss.
Algorithm 1 Computing the set $\mathcal{U}$ of feasible inclusion class pairs from the training set ground-truth annotations $\mathcal{S}$. The set $\mathcal{P}$ contains all pairs of non-identical class labels for the dataset taxonomy. We aim to check whether the first element of the pair, $p$, can include the second one, $q$. We are taking 5 pixels as the minimum threshold for a contour of $q$ completely overlapping with any segment of $p$. Function C returns (i) the number of connected segments of class $q$ in the input annotation, and (ii) a discrete image, with each pixel labeled with its connected component ID or 0 for the background. Function F fills the holes of the mask of the connected component at hand. Then, function Cont returns a binary image in which only pixels belonging to the outer contour of the respective connected component are set.
@@
4:$\texttt{count}=0$
@@
6:$\mathcal{U}\leftarrow\emptyset$
7:for  $(p,q)\leftarrow\mathcal{P}$ do
8:     for  $S\leftarrow\mathcal{S}$ do
9:         mask_1 $\leftarrow\mathbbm{1}[S=p]\in\mathbb{R}^{1\times H\times W}$
10:         mask_2 $\leftarrow\mathbbm{1}[S=q]\in\mathbb{R}^{1\times H\times W}$
11:         if mask_1 not empty $\land$ mask_2 not empty then
[…]
13:         end if
[…]
15:         for  $l\leftarrow 1$ to noOFLabels do
16:              comp_mask $\leftarrow(\mathbbm{1}[\texttt{component}=l])$
@@
20:              if $\textbf{NNZ}(\texttt{result})=\textbf{NNZ}(\texttt{contour})\land\textbf{NNZ}(\texttt{contour})\geq 5$ then
21:                  $\texttt{count}\leftarrow\texttt{count}+1$
22:                  break
23:              end if
24:         end for
25:     end for
[…]
27:         $\mathcal{U}\leftarrow\mathcal{U}\cup\{(p,q)\}$
28:         $\texttt{count}=0$
[…]
30:     end if
31:end for